# Advancing MAPF towards the Real World: A Scalable Multi-Agent Realistic Testbed (SMART)

### v1

@@ Abstract
We present Scalable Multi-Agent Realistic Testbed (SMART), a realistic and efficient software tool for evaluating Multi-Agent Path Finding (MAPF) algorithms. MAPF focuses on planning collision-free paths for a group of agents. While state-of-the-art MAPF algorithms can plan paths for hundreds of robots in seconds, they often rely on simplified robot models, making their real-world performance unclear. Researchers typically lack access to hundreds of physical robots in laboratory settings to evaluate the algorithms. Meanwhile, industrial professionals who lack expertise in MAPF require an easy-to-use simulator to efficiently test and understand the performance of MAPF algorithms in their specific settings. SMART fills this gap with several advantages: (1) SMART uses a physics-engine-based simulator to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF algorithms and robot models, and (3) SMART scales to thousands of robots. In addition, we use SMART to explore and demonstrate research questions about the execution of MAPF algorithms in real-world scenarios. The code is publicly available at https://jingtianyan.github.io/publication/2025-smart.
1 Introduction
Figure 1: (a) A simulation environment in SMART. (b) The original 2-D grid map from the MovingAI benchmark. (c) Robot model used in the environment. (d) A 3D view of the environment.
@@ 1 Introduction
The Multi-Agent Path Finding (MAPF) problem Stern et al. (2019) focuses on planning collision-free paths for large groups of robots within a known environment. State-of-the-art MAPF algorithms can plan paths for hundreds of robots within seconds Ma et al. (2019); Li et al. (2021c); Zhang et al. (2022); Li et al. (2022). However, these algorithms make several simplifying assumptions. First, they often rely on simplified robot models that ignore kinodynamic constraints while planning the robots’ paths. Second, they assume that robots can execute these paths perfectly, without accounting for uncertainties introduced by real-world factors. Third, they typically plan and evaluate paths on unrealistic discretized 2-D grid maps Stern et al. (2019). Therefore, while researchers have shown that MAPF algorithms have a wide range of applications Ma et al. (2017); Hönig et al. (2019); Ho et al. (2022); Li et al. (2023); Zhang et al. (2023b, c, 2024); Zang et al. (2025), the algorithms are mostly evaluated under simplified settings, leaving their performance in more realistic settings unclear.
@@
MAPF researchers in both academia and industry require an efficient, realistic, and scalable software tool to evaluate MAPF algorithms in realistic settings. In academia, deploying hundreds of physical robots in a large environment to evaluate MAPF algorithms is impractical. Meanwhile, developing software that can evaluate large groups of robots is a non-trivial task for MAPF researchers. Researchers must create high-fidelity simulation environments and robot models to capture realistic factors, as well as develop essential software components such as controllers and execution frameworks to execute their plans. Moreover, since many researchers come from the AI community, learning such simulation tools can be time-consuming and require significant effort. While some software tools Schaefer et al. (2023); Heuer et al. (2024) have been developed to evaluate MAPF algorithms in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF algorithms and maps. In addition, industrial professionals possess realistic simulators for their applications of interest. However, they may not have the resources and expertise to implement state-of-the-art MAPF algorithms in their simulators. By having a realistic simulator that can easily interface with existing implementations of MAPF algorithms, industrial professionals can more easily understand and select a MAPF algorithm that fits their demands.
@@
In this paper, we present SMART, a Scalable Multi-Agent Realistic Testbed, a software tool used to evaluate MAPF algorithms in realistic scenarios. SMART comprises three main components: (1) a physics-engine-based simulator based on Argos 3 Pinciroli et al. (2012), (2) an execution monitoring (EM) server, and (3) robot-specific executors. The simulator generates realistic simulation environments based on user-defined configurations. Fig. 1 shows an example environment in SMART created from a 2-D grid map from the MovingAI benchmark Stern et al. (2019). The EM server parses the user-provided MAPF paths, tracks the execution progress, and communicates with the executors. It leverages the Action Dependency Graph (ADG) Hönig et al. (2019), a MAPF execution framework, to ensure robust execution of the paths generated using various simplified robot models. Each robot is assigned an individual executor that receives and executes actions from the server. SMART incorporates key real-world factors, including robot kinodynamics, collision dynamics, communication delays, execution imperfections, and spatial inaccuracies caused by controller limitations.
Our contributions in this paper are: (1) We present SMART, an open-source software tool to evaluate MAPF algorithms in realistic scenarios, which can scale to thousands of robots while ensuring replicability, (2) we evaluate SMART with a variety of open-source implementations of MAPF algorithms and robot models, showing its flexibility, and (3) we show some important findings regarding the application of MAPF algorithms in realistic scenarios. We also use them as examples of how SMART can be used to address such problems.
@@
2 Background
2.1 MAPF Problem and MAPF Plan
@@
A MAPF instance Stern et al. (2019) consists of an undirected graph $\mathcal{G}_{M}=(\mathcal{V}_{M},\mathcal{E}_{M})$ and a set of $I$ robots =$\mathcal{R}=\{r_{1},\dots,r_{I}\}$ with their corresponding start and goal locations. Given a MAPF instance, a MAPF algorithm aims at finding a MAPF plan, consisting of collision-free paths for all robots from their start to goal locations. Standard MAPF uses a simplified robot model by discretizing the time into timesteps and assuming that, at each timestep, every robot either moves to an adjacent vertex or stays at its current vertex. Two robots collide if they are at the same vertex or swap vertices at the same timestep. An example MAPF instance and plan on a 2-D grid graph is shown in Fig. 2 (a) and (b), respectively. The objective of MAPF is to minimize the sum-of-cost, defined as the sum of travel time.
@@
Prior research has proposed many MAPF algorithms. Some methods focus on finding solutions with optimal or bounded-suboptimal guarantees Sharon et al. (2015); Barer et al. (2014); Li et al. (2021c), while others focus on scalability, solving MAPF problems involving thousands of robots Okumura et al. (2022); Okumura (2023). Anytime algorithms quickly find an initial solution and then iteratively refine it as time progresses, balancing solution quality with computational efficiency Li et al. (2021a); Huang et al. (2022). Additionally, some methods incorporate complex real-world factors, such as kinodynamics and unexpected delays, into the planning process Cohen et al. (2019); Atzmon et al. (2020); Yan and Li (2024a, b). They usually employ a two-level planner, where the high level employs MAPF-based methods to resolve collisions among agents, while the low level plans for each agent, consider real-world factors.
2.2 Executing MAPF Plans
@@
Given a MAPF plan, robots can remain collision-free if they strictly follow it spatially and temporally. While existing path-tracking techniques allow for reasonable spatial accuracy, precise temporal adherence is more challenging Čáp et al. (2016), especially since MAPF plans are typically generated using imperfect robot models. Thus, we need techniques to execute MAPF plans safely given the potential delays. One naive way is synchronized execution, where agents need to wait for all robots to finish their current action before starting the next. This ensures strict compliance with the MAPF plan but significantly reduces efficiency. Alternatively, one can replan whenever a robot is delayed, but this is computationally heavy and thus does not scale well. Therefore, executing MAPF plans using an execution framework is preferred Čáp et al. (2016). The state-of-the-art execution framework is Action Dependency Graph (ADG) Hönig et al. (2019), which enforces passing orders defined by MAPF plans while accounting for delays. ADG can execute plans generated by MAPF planners of various simplified robot models Varambally et al. (2022).
@@
An ADG, as shown in Fig. 2 (c), is a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{1},\mathcal{E}_{2})$ that encodes dependencies between actions from a MAPF plan. The vertex set $\mathcal{V}=\{v_{i}^{k}:k\in[1,n^{i}],r_{i}\in\mathcal{R}\}$ represents all actions performed by each robot $r_{i}$, where each vertex $v_{i}^{k}$ corresponds to the $k$-th action of $r_{i}$, and $n^{i}$ is the number of actions of $r_{i}$. The edge set $\mathcal{E}_{1}$ includes Type-1 edges, which enforces the sequential execution of actions for each robot. For a robot $r_{i}$, a Type-1 edge $(v_{i}^{k},v_{i}^{k+1})$ ensures that action $v_{i}^{k}$ is completed before action $v_{i}^{k+1}$ starts. The edge set $\mathcal{E}_{2}$ includes Type-2 edges, which enforces the inter-robot passing order at shared vertices. For any two vertices $v_{i}^{k}$ and $v_{j}^{s}$ from $r_{i}$ and $r_{j}$, a Type-2 edge $(v_{i}^{k},v_{j}^{s})$ ensures that $v_{i}^{k}$ must be completed before $v_{j}^{s}$ can start. During execution, each action is performed according to these dependencies. The detailed execution procedure is described in Section 3.2.
2.3 Evaluating MAPF Plans
@@
To evaluate MAPF algorithms, researchers typically generate a plan for a given MAPF instance and then execute it. While many MAPF problem benchmark instances are proposed for discretized 2-D-grid maps Stern et al. (2019); Gebser et al. (2018); Sturtevant (2012); Mohanty et al. (2020); Qian et al. (2024), few software tools are provided to execute the MAPF plans in more realistic environments. Therefore, most prior works are evaluated using simplified robot models that assume no delays and discretized time and space.
@@
Meanwhile, some software tools are presented to evaluate the MAPF plans in more realistic settings. MRP-Bench Schaefer et al. (2023) leverages a high-fidelity, physics-based simulator Gazebo Koenig and Howard (2004) to simulate multi-robot systems. It integrates low-level controllers to account for robots’ kinodynamic constraints and potential delays. REMROC Heuer et al. (2024) also uses Gazebo to realistically evaluate MAPF plans, with a particular focus on human-shared environments. However, they have two main limitations. First, they often require a lot of engineering effort to use user-defined maps and MAPF planners, limiting their flexibility for different scenarios. Second, their scalability is constrained by the high computational overhead of Gazebo, which models many factors less relevant to MAPF research. MRP-Bench, for example, does not scale to more than 10 robots. Additionally, many MAPF researchers from the AI community may not be familiar with tools like Gazebo, making installation and usage a challenging process.
3 SMART Architecture
@@
3.1 Simulator
[…]
@@ 3.1 Simulator
3.2 Execution Monitoring (EM) Server
@@
During the initialization, we use the MAPF plan parser to construct an ADG. We pre-define the format to store the MAPF plan as a sequence of timed locations for each robot. The parser can take as input any valid MAPF plans of this format, regardless of the MAPF planners that generate the plans. This allows SMART to work with plan using varying simplified robot models. The parser first converts the path of each robot into a sequence of actions. Specifically, the parser infers the orientation of a robot by comparing consecutive vertices along its path. If the required orientation differs from the current orientation, a rotate action is added. The ADG is constructed based on the sequence of actions. ADG requires the MAPF plan to be 1-robust, ensuring that robots do not collide even if delayed by one timestep. However, ensuring a 1-robust plan is not standard in most MAPF models. As a result, if we directly build an ADG based on a non-1-robust plan, we may create deadlocks, such as those illustrated in Fig. 4 (a). While these robots are planned to rotate simultaneously in a cycle in the non-1-robust plan, with ADG, they will end up waiting indefinitely for others to complete their next actions. To accommodate a broader range of planners, motivated by Hönig et al. (2016), we add an extra vertex between consecutive vertices in SMART as shown in Fig. 4 (b), making the ADG compatible with these MAPF models.
@@
During execution, an execution monitor framework manages the progress of execution and communicates with the executors. This framework follows the method described in ADG Hönig et al. (2019) to track and manage execution. Each vertex in the ADG represents an action and can have one of three statuses: staged, enqueued, or finished. At first, all vertices are staged. A vertex transitions to the enqueued if it has no preceding vertices or (1) its preceding vertex connected by a Type 1 edge is either enqueued or finished, and (2) its preceding vertices connected by Type 2 edges are finished. Actions associated with enqueued vertices are then sent to the executors for execution, while other vertices remain staged. A vertex is marked as finished when the executor confirms the completion of its associated action with the server. Communication with the executors is managed by an RPC server built with rpclib [^1] Available at http://rpclib.net..
3.3 Executor

### v2

@@ Abstract
We present Scalable Multi-Agent Realistic Testbed (SMART), a realistic and efficient software tool for evaluating Multi-Agent Path Finding (MAPF) algorithms. MAPF focuses on planning collision-free paths for a group of agents. While state-of-the-art MAPF algorithms can plan paths for hundreds of robots in seconds, they often rely on simplified robot models, making their real-world performance unclear. Researchers typically lack access to hundreds of physical robots in laboratory settings to evaluate the algorithms. Meanwhile, industrial professionals who lack expertise in MAPF require an easy-to-use simulator to efficiently test and understand the performance of MAPF algorithms in their specific settings. SMART fills this gap with several advantages: (1) SMART uses physics-engine-based simulators to create realistic simulation environments, accounting for complex real-world factors such as robot kinodynamics and execution uncertainties, (2) SMART uses an execution monitor framework based on the Action Dependency Graph, facilitating seamless integration with various MAPF algorithms and robot models, and (3) SMART scales to thousands of robots. The code is publicly available at https://github.com/smart-mapf/smart.
Index Terms:
Multi-Agent Path Finding, Multi-Robot System, Realistic Simulation.
I Introduction
@@
The Multi-Agent Path Finding (MAPF) problem [1] focuses on planning collision-free paths for a large group of robots within a known environment. This problem finds various real-world applications [2, 3, 4], with warehouse automation being a prominent example, where hundreds of robots must be coordinated to perform daily operations. In recent years, MAPF has grown into a rapidly expanding research area [5, 6]. State-of-the-art MAPF algorithms can plan paths for hundreds of robots within seconds [7, 8, 9]. However, these algorithms make several simplifying assumptions. First, they rely on simplified robot models that ignore kinodynamic constraints while planning the robots’ paths. Second, they assume that robots can execute these paths perfectly, without accounting for uncertainties introduced by real-world factors. Therefore, while researchers have shown that MAPF algorithms have a wide range of applications, their performance in more realistic settings is unclear.
Figure 1: (a) A simulation environment in SMART. (b) The original 2-D grid map from the MovingAI benchmark. (c, d) Robot models used in the environment. (e) A 3D view of the environment.
@@
MAPF researchers in both academia and industry require an efficient, realistic, and scalable software tool to evaluate MAPF algorithms in realistic settings. In academia, deploying hundreds of physical robots in a large environment to evaluate MAPF algorithms is impractical. Meanwhile, developing software that can evaluate large groups of robots is a non-trivial task for MAPF researchers. Researchers must create high-fidelity simulation environments and robot models to capture realistic factors, as well as develop essential software components such as controllers and execution frameworks to execute their plans. Moreover, since many researchers come from the AI community, learning such simulation tools can be time-consuming and require significant effort. While some software tools [10, 11] have been developed to evaluate MAPF algorithms in realistic settings, they do not scale to hundreds of robots and require significant engineering efforts to integrate new MAPF algorithms and maps. In addition, industrial professionals possess realistic simulators for their applications of interest. However, they may not have the resources and expertise to implement state-of-the-art MAPF algorithms in their simulators. By having a realistic simulator that can easily interface with existing implementations of MAPF algorithms, industrial professionals can more easily understand and select a MAPF algorithm that fits their demands.
@@
In this paper, we present SMART, a Scalable Multi-Agent Realistic Testbed, a software tool used to evaluate MAPF algorithms in realistic scenarios. SMART comprises three main components: (1) physics-engine-based simulators, (2) an execution monitoring (EM) server, and (3) robot-specific executors. The simulator generates realistic simulation environments based on user-defined configurations. Fig. 1 shows an example environment in SMART created from a 2-D grid map from the MovingAI benchmark [1]. The EM server parses the user-provided MAPF paths, tracks the execution progress, and communicates with the executors. It leverages the Action Dependency Graph (ADG) [3], a MAPF execution framework, to ensure robust execution of the paths generated using various simplified robot models. Each robot is assigned an individual executor that receives and executes actions from the server. SMART incorporates key real-world factors, including robot kinodynamics, collision dynamics, communication delays, execution imperfections, and temporal uncertainty during execution.
Our contributions in this paper are as follows: 1. We present SMART, an open-source software tool for evaluating MAPF algorithms in realistic scenarios. SMART directly accepts MAPF plans and problem instances from existing benchmarks. 2. We integrate SMART with two simulators, ARGoS3 and Isaac Sim, and robot models and empirically evaluate its scalability and replicability. SMART can scale to thousands of robots while maintaining consistent and reproducible results, while prior tools fail to handle a hundred robots. 3. We develop an online interface with interactive visualization, allowing users to easily monitor execution progress and access detailed statistical data during simulation. 4. We validate SMART on real-world mobile robots, demonstrating its robustness and effectiveness in physical environments.
@@
II Background
In this section, we first introduce the MAPF problem and related work. Next, we discuss existing execution frameworks for executing MAPF plans. Finally, we review benchmarks used to evaluate MAPF algorithms.
II-A MAPF Problem and MAPF Plan
@@
A MAPF instance [1] consists of an undirected graph $\mathcal{G}_{M}=(\mathcal{V}_{M},\mathcal{E}_{M})$ constructed from a 2-D grid map and a set of $I$ robots $\mathcal{R}=\{r_{1},\dots,r_{I}\}$ with their corresponding start and goal locations. Given a MAPF instance, a MAPF algorithm aims at finding a MAPF plan, consisting of collision-free paths for all robots from their start to goal locations. Standard MAPF uses a simplified robot model by discretizing the time into timesteps and assuming that, at each timestep, every robot either moves to an adjacent vertex or stays at its current vertex. Two robots collide if they are at the same vertex or swap vertices at the same timestep. An example MAPF instance and plan on a 2-D grid graph are shown in Fig. 2 (a) and (b), respectively. The objective of MAPF is to minimize the sum-of-cost, defined as the sum of travel time.
@@
Prior research has proposed many MAPF algorithms, most of which are developed based on 2-D grid maps. Some methods focus on finding solutions with optimal or bounded-suboptimal guarantees [12, 13], while others focus on scalability, solving MAPF problems involving thousands of robots [14, 15]. Anytime algorithms quickly find an initial solution and then iteratively refine it as time progresses, balancing solution quality with computational efficiency [16, 17]. Additionally, some methods incorporate complex real-world factors, such as kinodynamics and unexpected delays, into the planning process [18, 19, 20, 21]. They usually employ a two-level planner, where the high level employs MAPF-based methods to resolve collisions among agents, while the low level plans for each agent, consider real-world factors.
II-B Executing MAPF Plans
@@
Given a MAPF plan, robots can remain collision-free if they strictly follow it spatially and temporally. While existing path-tracking techniques allow for reasonable spatial accuracy, precise temporal adherence is more challenging [22], especially since MAPF plans are typically generated using imperfect robot models. Thus, we need techniques to execute MAPF plans safely given the uncertainty in execution time. One naive way is synchronized execution, where agents need to wait for all robots to finish their current action before starting the next. This ensures strict compliance with the MAPF plan but significantly reduces efficiency. Alternatively, one can replan whenever a robot is delayed, but this is computationally heavy and thus does not scale well. Therefore, executing MAPF plans using an execution framework is preferred [22]. The state-of-the-art execution framework is Action Dependency Graph (ADG) [3], which enforces passing orders defined by MAPF plans while accounting for delays. ADG can execute plans generated by MAPF planners of various robot models [23].
@@
An ADG, as shown in Fig. 2 (c), is a directed graph $\mathcal{G}=(\mathcal{V},\mathcal{E}_{1},\mathcal{E}_{2})$ that encodes dependencies between actions from a MAPF plan. The vertex set $\mathcal{V}=\{v_{i}^{k}:k\in[1,n^{i}],r_{i}\in\mathcal{R}\}$ represents all actions performed by each robot $r_{i}$, where each vertex $v_{i}^{k}$ corresponds to the $k$-th action of $r_{i}$, and $n^{i}$ is its number of actions. The edge set $\mathcal{E}_{1}$ includes Type-1 edges, which enforce the sequential execution of actions by each robot. For a robot $r_{i}$, a Type-1 edge $(v_{i}^{k},v_{i}^{k+1})$ ensures that action $v_{i}^{k}$ is completed before action $v_{i}^{k+1}$ starts. The edge set $\mathcal{E}_{2}$ includes Type-2 edges, which enforce the inter-robot passing order at shared vertices. For any two vertices $v_{i}^{k}$ and $v_{j}^{s}$ from $r_{i}$ and $r_{j}$, a Type-2 edge $(v_{i}^{k},v_{j}^{s})$ ensures that $v_{i}^{k}$ must be completed before $v_{j}^{s}$ can start. During execution, each action is performed according to these dependencies. The detailed execution procedure is described in Section III-B.
II-C Evaluating MAPF Plans
@@
To evaluate MAPF algorithms, researchers typically generate a plan for a given MAPF instance and then execute it. While many MAPF benchmarks with 2-D-grid maps are proposed [1, 24, 25, 26], most prior MAPF works are evaluated on them using simplified robot models that assume no delays and discretized time and space. Some benchmarks have attempted to include realistic elements such as acceleration and rotation [27, 28]. However, their model is still far from realistic, and typically requires the same robot models for both planning and execution phases.
@@
Given the difficulty of modeling all real-world factors analytically, physics-engine-based simulators offer a promising approach for more accurate evaluation. MRP-Bench [10] leverages a high-fidelity, physics-based simulator Gazebo [29] to simulate multi-robot systems. It integrates low-level controllers to account for robots’ kinodynamic constraints and potential delays. REMROC [11] also uses Gazebo to realistically evaluate MAPF plans, with a particular focus on human-shared environments. However, these existing tools have two main limitations. First, they often require significant engineering effort to integrate user-defined maps and MAPF planners, limiting their flexibility for different scenarios. In particular, they lack a general execution framework, making it difficult to directly evaluate MAPF planners using different robot models. Second, they have limited scalability: the largest number of robots tested in previous work is less than 100 agents. Most of these works rely on Gazebo, whose high computational overhead, resulting from modeling many factors not relevant to MAPF, significantly constrains scalability. MRP-Bench, for example, does not scale to more than 10 robots.
III SMART Architecture
@@
III-A Simulator
Given a user-provided configuration, the simulator creates realistic simulation environments. The configuration specifies obstacle locations and shapes, robot start locations, and detailed robot configurations (e.g., the robot model, as well as its velocity and acceleration limits). To support the widely used MovingAI benchmark [1], SMART can converts its 2-D grid maps and corresponding MAPF instances into the configuration required by the simulator. During execution, the simulator uses a physics engine to update the states of the robots and environment, including their poses, dynamics, sensors, and actuators, based on the received control commands from the executors. The executors can query the simulator for real-time robot information, such as their current pose. To accommodate diverse user requirements, SMART supports two simulation platforms: ARGoS3 [30] and Isaac Sim [31].
III-A1 ARGoS3
ARGoS3 is a multi-robot simulator with excellent scalability, advanced physics engines, and support for custom robot and environment models, making it suitable for our tasks. While users can specify any robot model, we use the foot-bot (shown in Fig. 1 (c)), a differential-drive robot, in our experiments. It closely resembles robots used in real-world warehouses and can move forward and rotate in place.
III-A2 Isaac Sim
Isaac Sim is a high-fidelity robotics simulator developed by NVIDIA, featuring GPU-accelerated physics, photorealistic rendering, and seamless integration with ROS. It is well-suited for sim-to-real transfer and physically grounded perception and control tasks. We use the JetBot robot (shown in Fig. 1 (d)), a differential-drive robot commonly used in research and education.
@@ III-A2 Isaac Sim
III-B Execution Monitoring (EM) Server
@@
During the initialization, we use the MAPF plan parser to construct an ADG. We define a standardized format to store the MAPF plan as a sequence of timed locations for each robot $r_{i}$, i.e., $((x_{i}^{0},y_{i}^{0},t_{i}^{0}),...,(x_{i}^{T},y_{i}^{T},t_{i}^{T}))$. The parser can take as input any valid MAPF plans of this format, regardless of the MAPF planners that generate the plans. This allows SMART to work with plans using varying simplified robot models. The parser first converts the path of each robot into a sequence of actions. Specifically, the parser infers the orientation of a robot by comparing consecutive vertices along its path and adds a rotation action when needed. The ADG is constructed based on the sequence of actions. ADG requires the MAPF plan to be circle-conflict free, ensuring that robots do not collide even if delayed by one timestep [3]. However, ensuring this is not standard in most MAPF models. As a result, if we directly build an ADG based on a plan with circle conflict, we may create deadlocks, such as the one illustrated in Fig. 4 (a). While these robots are planned to rotate simultaneously in a cycle, with ADG, they will end up waiting indefinitely for others to complete their next actions. To accommodate a broader range of planners, motivated by [32], we add an extra vertex between consecutive vertices in SMART as shown in Fig. 4 (b), making the ADG compatible with these MAPF models.
@@
During execution, an execution monitor framework manages the progress of execution and communicates with robot executors. This framework follows the method described in ADG [3] to track and manage execution. Each vertex in the ADG represents an action and can have one of three statuses: staged, enqueued, or finished. At first, all vertices are staged. A vertex transitions to enqueued if (1) it has no preceding vertices, or (2) its preceding vertex connected by a Type-1 edge is enqueued or finished, and all its preceding vertices connected by Type-2 edges are finished. Actions associated with enqueued vertices are then sent to the executors for execution. A vertex is marked as finished when the executor confirms the completion of its associated action with the server. Communication with robot executors is managed by an RPC server built with rpclib [^1] Available at http://rpclib.net..
III-C Executor